# thanks: corresponding author

### v1

@@ 4.2 Entanglement entropy
Figure 10: (a) $\overline{\langle\mathcal{E}\rangle}$ versus $N$ for several values of $\gamma$ and for fixed $h=1$, $K=2$. The curves refer to different values of $\gamma_{j}=j\,\delta\gamma$ with $\delta\gamma=0.1$, from $j=1$ (darkest) to $j=10$ (lightest). The largest Lyapunov exponent $\lambda$ is positive for $\gamma\lesssim 0.87$ (curves with circles) and negative otherwise (curves with squares). (b) Position $N^{*}$ of the maximum of $\overline{\langle\mathcal{E}\rangle}$ in panel (a) versus $\gamma$, log-log scale. (c) Maximum $\overline{\langle\mathcal{E}\rangle}$ divided by $N^{*}$ versus $\gamma$, linear scale.
In the central and lower rows of Fig. 8, we show the EE versus time in the presence of dissipation, being it either weak ($\gamma=0.2$, middle row) or strong ($\gamma=0.5$, bottom row), for the same values of $K$ considered before. Panels (c),(e), and (f) correspond to a regular mean-field dynamics ($\lambda<0$), while panel (d) is for a chaotic dynamics ($\lambda>0$). We observe that, after a transient, all curves converge to an asymptotic value. Further insight can be obtained by focusing on the dependence of the asymptotic value $\overline{\langle\mathcal{E}\rangle}$ on $K$ and on the system size $N$.
In Fig. 9 we show the behavior of $\overline{\langle\mathcal{E}\rangle}$ versus $K$, for the same values of $\gamma$ and system sizes used in Fig. 8. First of all, we notice a weak relation between such curves and the sign of the largest Lyapunov exponent $\lambda$ (see the color code in the background, as in Fig. 5). Furthermore, for increasing $N$, the values of $\overline{\langle\mathcal{E}\rangle}$ display a tendency to saturate to a finite limit (area-law behavior) and even to decrease. To better understand this point, in Fig. 10 panel (a) we report the asymptotic value $\overline{\langle\mathcal{E}\rangle}$ versus $N$, for fixed $K$ and different values of $\gamma$. As a result of the competition between driving and dissipation, we observe a nonmonotonic behavior with the size $N$. We fix $K$ and $h$ because, with this choice of parameters, the maximum can be observed more easily, but this is a general phenomenon. In fact, the nonmonotonicity occurs both for an underlying classical chaotic (curves with circles) and a regular (curves with squares) regime. However, our data suggest that all the curves eventually follow an area-law behavior, where $\overline{\langle\mathcal{E}\rangle}$ becomes independent of $N$.
@@ 4.3 Magnetization fluctuations
Figure 11: Variance of $m_{z}$ over time and randomness [Eq. (22)] as a function of the size, up to $N=640$. The largest Lyapunov exponent is positive for $\gamma\lesssim 0.5$ (curves with circles) and negative otherwise (curves with squares), as shown in Fig. 2a. We fix $K=5$ and $h=0.5$.

### v3

@@ 4.2 Entanglement entropy
Figure 6: (a) $\overline{\langle\mathcal{E}\rangle}$ versus $N$ for several values of $\gamma$ and for fixed $h=1$, $K=2$. The curves refer to different values of $\gamma_{j}=j\,\delta\gamma$ with $\delta\gamma=0.1$, from $j=1$ (darkest) to $j=10$ (lightest). The largest Lyapunov exponent $\lambda$ is positive for $\gamma\lesssim 0.87$ (curves with filled circles) and negative otherwise (curves with empty squares). (b) Position $N^{*}$ of the maximum of $\overline{\langle\mathcal{E}\rangle}$ in panel (a) versus $\gamma$, log-log scale. (c) Maximum $\overline{\langle\mathcal{E}\rangle}$ divided by $N^{*}$ versus $\gamma$, linear scale.
As done in the previous section, we studied the entanglement dynamics, averaged over the trajectories, as a function of the stroboscopic time $t_{n}$. In all cases, we see that the entropy increases and then saturates to an asymptotic value (see Appendix B). Further insight can be obtained by focusing on the dependence of the asymptotic value $\overline{\langle\mathcal{E}\rangle}$ on $K$ and on the system size $N$. This is reported in Appendix B, where we show that, by increasing $N$, the values of $\overline{\langle\mathcal{E}\rangle}$ display a tendency to saturate to a finite limit (area-law behavior) and even to decrease. To better understand this point, in Fig. 6 panel (a) we report $\overline{\langle\mathcal{E}\rangle}$ versus $N$, for fixed $K$ and different values of $\gamma$. As a result of the competition between driving and dissipation, we observe a nonmonotonic behavior with the size $N$. We fix $K$ and $h$ because, with this choice of parameters, the maximum can be observed more easily, but this is a general phenomenon. In fact, the nonmonotonicity occurs both for an underlying classical chaotic (curves with circles) and a regular (curves with squares) regime. However, our data suggest that all the curves eventually follow an area-law behavior, where $\overline{\langle\mathcal{E}\rangle}$ becomes independent of $N$.
@@ 4.3 Magnetization fluctuations
Figure 7: Variance of $m_{z}$ over time and randomness [Eq. (22)] as a function of the size, up to $N=640$. The largest Lyapunov exponent is positive for $\gamma\lesssim 0.5$ (curves with filled circles) and negative otherwise (curves with empty squares), as shown in Fig. 2(a). We fix $K=5$ and $h=0.5$.
@@ A.2 Hausdorff dimension
[…]
Appendix B Additional numerical results
In this section, we provide additional numerical results concerning nonstabilizerness and entanglement dynamics to accompany the results discussed in the main text.
B.1 Nonstabilizerness
Figure 12 reports the behavior of the nonstabilizerness $\langle\mathcal{M}_{2}(t_{n})\rangle$ versus the stroboscopic time $t_{n}$ in the dissipative quantum kicked top, after averaging over the randomness of the quantum trajectories. We first notice that the SRE as a function of time rapidly tends to an asymptotic value. Moreover, when the kick strength $K$ is small and the dynamics is regular, the asymptotic value seems to grow extensively with $N$ (left panels, for $K=0$). On the other hand, for larger kick strengths, the dynamics becomes chaotic and the asymptotic value increases more slowly, possibly as $\ln N$ (right panels, for $K=5$).
To be more quantitative, we first extrapolate the asymptotic value of the SRE as discussed in Sec. 4. We plot it versus the system size (up to $N=256$) in Fig. 13, for various combinations of the kick strength $K$ and the decay rate $\gamma$, corresponding to different values of the Lyapunov exponents of the undarlying semiclassical dynamics [see Fig. 2(a)]. We observe that, for parameters that correspond to regular semiclassical dynamics ($\lambda\leq 0$), the asymptotic SRE is extensive and grows as $\overline{\langle\mathcal{M}_{2}\rangle}\sim N^{\alpha}$, with $\alpha$ close to one. On the other hand, in the chaotic regime, the SRE grows more slowly. Indeed, a numerical fitting returns an exponent that is appreciably smaller than one (see Tab. 1). Given the relatively small available sizes, by simply analyzing these data, we cannot rule out the possibility that the scaling is sub-extensive (but still power law), rather than logarithmic, which is why, in the main text, we focus on the more interesting scaling of $\delta\mathcal{M}_{N}$, see Sec. 4.1.
B.2 Entanglement entropy
In the central and lower rows of Fig. 14, we show the EE versus time in the presence of dissipation, being it either weak ($\gamma=0.2$, middle row) or strong ($\gamma=0.5$, bottom row), for the same values of $K$ considered before. Panels (c),(e), and (f) correspond to a regular mean-field dynamics ($\lambda<0$), while panel (d) is for a chaotic dynamics ($\lambda>0$). We observe that, after a transient, all curves converge to an asymptotic value. Further insight can be obtained by focusing on the dependence of the asymptotic value $\overline{\langle\mathcal{E}\rangle}$ on $K$.
In Fig. 15 we show the behavior of $\overline{\langle\mathcal{E}\rangle}$ versus $K$, for the same values of $\gamma$ and system sizes used in Fig. 14. First of all, we notice a weak relation between such curves and the sign of the largest Lyapunov exponent $\lambda$ (see the color code in the background, as in Fig. 3). Furthermore, for increasing $N$, the values of $\overline{\langle\mathcal{E}\rangle}$ display a tendency to saturate to a finite limit (area-law behavior) and even to decrease.
In the main text, we focus our attention on the scaling of the asymptotic value of the EE with $N$.